OpenStudy (anonymous):

(Sinx+cosx)(1-sinxcosx) can be proved to be made sin^3x+cos^3x. How do I do this? I can't simplify the sinxcosx part

OpenStudy (raden):

use this formula : a^3 + b^3 = (a+b)(a^2 - ab + b^2) for this case, use a = sinx and b = cosx